68. A force F a Fi Fyj with F 50 N and F 12 N is exerted on a particle as the particle moves along the x axis from x 1.0 m to x 5.0 m. (a) Determine the work done by the force on the particle. (b) What is the angle be- tween the force and the particles displacement?
Answer #1

(b)Angle: ta12 F 50 12 50 -1 θ=13.5° F-+F-50 +12-5142N d=-5-1=-6m Workdone: W-Fa cos θ= 51 .42x_6 xcos 13.5°ー-298.71 J W-298.